OpenStudy (anonymous):

Get a fan and a medal plz help me! A painter used 5/6 gallon of paint for one job and 1/4 gallon for another job. How many gallons of paint did the painter use for both jobs? A.1/2gal B.3/5gal C.1 and 1/12gal D.1 and 1/16gal Already solved

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@Imad_Sawal

OpenStudy (anonymous):

The total amount of Paint : (5 / 6) + (1 / 4 ) Taking the L.C.M : = ( (5 * 4 ) + ( 1 * 6 ) ) / 24 = 20 + 6 / 6 \[ 1 \frac{ 1 }{ 12 }\]He used 1 1/12 gallons of paint in total !

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

you think you can help me with another?

OpenStudy (anonymous):

The base of a triangle is 5/8 ft. The height is 1/3 ft. How much greater is the base of the triangle than the height? Express your answer in lowest terms.

OpenStudy (anonymous):

ok i know i have to find the same denominator so the factors of 8 is 8 16 24 32 40 and the factors of 3 is 3 6 9 12 15 18 21 24 i learned that 24 is the same denominator i got to that point but now i got lost after that.

OpenStudy (anonymous):

One basic mistake : 3,6,9,12 ... aren't the "Factors" of 3 , they are its "Multiples" 3 is a "factor" of 3,6,9,12 ... . http://www.mathsisfun.com/numbers/factors-multiples.html

OpenStudy (anonymous):

ok so that was a small mistake

OpenStudy (anonymous):

This is what I got. 5/8=15/24 1/3=8/24 than i subtract 15/24 and 8/24= 7/24

OpenStudy (anonymous):

 What is the question asking for ? How much is the "base" greater than the height ? In other words tell me by which length is the base more than the height. So the answer will be : base - height = ( 5 / 8 ) - ( 1/3 ) *Taking L.C.M* = (5*3) - (1*8) / 24 = 7 / 24 ft. ( The "Units" that is "Feet" are very important i.e can make you loose marks ) Right, there. I looked over your recent update just now .

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I have one more question

OpenStudy (anonymous):

Abby walked 3/4mile on Saturday and 7/12mile on Sunday. How many miles did she walk in all? A. 5/8mi B. 5/6mi C. 1 and 1/3 D. 1 and 1/2

OpenStudy (anonymous):

@sleepyjess